Given the equation of the circle x^2+y^2=225,Find
dy/dx
The equation of the tangent to the circle x^2+y^2=225 at the point (3,4)
1
Expert's answer
2016-03-09T08:36:42-0500
2xdx + 2y dy = 0, from here dy/dx = - x/y. equation of the tangent line y = dy/dx(x_0)(x - x_0) + y(x_0) = -3/4(x - 3) + 4 = -3/4x + 19/4